OpenStudy (anonymous):

The points (−7,10) and (−r,1) lie on a line with slope 3/4. Find the missing coordinate r.

OpenStudy (anonymous):

use slope formula: \[\frac{ y_2-y_1 }{ x_2 -x_1}=slope\]

OpenStudy (alexandervonhumboldt2):

now substitute your points and solve:\[\frac{ 10-1 }{-7+r }=3/4\]

OpenStudy (alexandervonhumboldt2):

solve for r